OpenStudy (anonymous):

I don't understand one part of this question please help. A cork has a volume of 4cm^3 and weights 10^-3N. What is the specific gravity of cork? m=F/g=10^-2/10=10^-3 kg Density=m/v (10^-3/4cm^3)*(10^2cm/1m)^3=250kg/m^3 Where did (10^2cm/1m)^3 came from? Please help this remains a great mystery to me

OpenStudy (chmvijay):

LOL u have to convert he NEWTON unit to gram :) 10^-3 N = 0.1 gram

OpenStudy (unklerhaukus):

\[10^2[\text {cm}]=100[\text {cm}]=1[\text m]\] \[\implies\frac{10^2\text {cm}}{1[\text m]}=1\] \[\implies\left(\frac{10^2\text {cm}}{1[\text m]}\right)^3=1^3=1\] this has been used to convert cubic centimeters to cubic meters

OpenStudy (chmvijay):

and also volume 4cm^3 to 4*10^-2 m^3

OpenStudy (chmvijay):

density = mass/ volume = 0.1 g /4 *10^-2 = 2.5 g/m3 =

OpenStudy (chmvijay):

hey is that right or what you did is right :?????